OpenStudy (anonymous):

jack mows his lawn in 4 hours. Marilyn mows her lawn in 5 hours. how long would it take them to do it together?

OpenStudy (anonymous):

1/5+1/4=1/x 4x+5x=20 x=20/9

hero (hero):

\[\frac{4 \times 5}{4+5}\]